SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (, 2007)

Attendees: Debbie Bard, Rainer Bartoldus, Brandon Drummond, Al Eisner, Fang Fang, Kim Hojeong, Su Dong, Mark Tibbetts.

Operations

  • Since the ROD last week, GLT transmission errors have started appearing (see JAS screenshots here and here). It does not indicate a real hardware problem, as that would not show up over all channels. It's more likely a buffer problem, possibly due to burstiness of trickle injection overloading the buffers. Can check this by re-making the plots offline and applying the physics window, which will remove events in the trickle window. This also ocurred at the end of Run 5, but the puzzler is why we didn't see any errors until now in Run 6? Check with the dataflow people to see if they did any work during the ROD that may have triggered this. To fix: the easiest thing to do would be to change the GLT firmware to back off the DAQ delay, by ~10 microseconds, provided it doesn't affect the deadtime. In any case, JAS plots could be updated to re-define the errors as a % of L1 accepts, and action (warning? page? flag run?) decided accordingly.

EMT

  • Still to fix damaged TRBs.
  • See last week's plots.
    Plot 1a shows EMT "energy" - pure ADC counts x0.4 with no selections on events. Errors are indeed binomial. Differences could be related to backgrounds in the two runs, or EMT timing.
    Plot 3a compares EMT "energy" (red) and EMC energy (black) for the same run. The difference shows how bad the 0.4 conversion factor is.

DCT/DCZ

  • Attempted to take a DCH-DCZ teststand calibration last week and ran into an number of problems. First the DCH ROM had a bad network problems, but even after that was fixed the calibration showed DCH hits in the xtc file but no TSF hits. The DCH ought to be sending out GLink frames every 60MHz but it appears to be sending 15 or 30MHZ. The TIOM has been replaced (thanks Karl!) and we'll try again this week.

L3

  • Tracking improvement - an xtc file with full statistics has been produced for the tracking group to study. Rainer found a possible problem with the geometry file while uploading it to the database, implying that the geometry used by the new tracking algorithm is transformed twice (the tranformation interpolates the radii at endplate to the radii at the centre of the wire). However, if this were the case the tracking should perform worse, not better! Needs looking into.
  • The DHP disconnect from server problem is on-going and occurs aout once a week still, and is often missed by the shifters. Jim is going to make it an alarm in ORC that cannot be ignored. The problem of the biased L3 lumi remains - could perhaps be fixed by using an average of all 30 nodes rather than the output?

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.